If liquid is combined in a 3 gallon to 1 gallon mix ratio how many batches would have to make if need 18 gallons
Answer by stanbon(75887)

If liquid is combined in a 3 gallon to 1 gallon mix ratio how many batches would have to make if need 18 gallons
--------------
3x + x = 18
4x = 18
x = 9/2
------
Amt. of 3 gallon liquid:: 3x = 3(9/2) = 27/2 = 13.5 gallons
Amt. of 1 gallon liquid:: x = 9/2 gallons
--------------
Cheers,
Stan H.
-----------